Solve the following 2 x 2 systems using substitutions
y= 2x +2
y= 1/2x + 8

Answer by elima(1433)

y= 2x +2
y= 1/2x + 8
Substitute 2x+2 for y in second equation;
2x+2 = 1/2x+8
2x-1/2x=8-2
3/2x=6
x=12/3=4
Now replace x with 4;
y = 2(4)+2
y=8+2=10
answer y = 10, x=4
:)